OpenStudy (samigupta8):

find the equation to the straight line which bisects and is perpendicular to the straight line joining the points A(a,b) and B(a',b')

hartnn (hartnn):

Can you find the slope of this line ? which joins the points A(a,b) and B(a',b') ?

hartnn (hartnn):

and since the given point bisects this line, also find the midpoint

OpenStudy (samigupta8):

it wil be b'-b /a'-a

OpenStudy (samigupta8):

i hve the equation of AB as y-b=\[\frac{ b'-b }{ a'-a }(x-a)\]

hartnn (hartnn):

yes, correct. and you know what is the product of slopes of perpendicular lines ?

OpenStudy (samigupta8):

\[i.e. y(a'-a)-x(b'-b)=a'b-ab'\]

OpenStudy (samigupta8):

equation to line perpendicular to AB is of form \[(b'-b)y+(a'-a)x+k=0\]

hartnn (hartnn):

correct, so u only need k now, right ?

hartnn (hartnn):

since the required line bisects AB, if we find the mid-point, that point will lie on the required line, right ? got this ?

OpenStudy (samigupta8):

yaa how am i going to get its value

hartnn (hartnn):